Consider two people being randomly selected. (For simplicity, ignore leap years.) (a) What is the probability that two people born in November have a birthday in the first half of the month? answer: (b) What is the probability that two people were born on Sunday? answer:
Answer #1

Answer:

a)

Given,

To determine the probability that two people born in november have a birthday in the first half of the month

Here in a year we have 365 days

i.e.,

n(S) = 365

And in a november month, we have 30 days, so the first half of month have 15 days

i.e.,

n(A) = n(B) = 15

Now consider,

P(A) = n(A)/n(S)

substitute values

P(A) = 15/365

Here A & B are independent events

Now the required probability is given as follows

i.e.,

P(A B) = P(A) * P(B)

= 15/365 *15*365

= 0.0017